OpenStudy (taylovesmatthew):

Find the measure of the central angle that you would draw to represent 47% in a circle graph. Round your answer to the nearest degree.

OpenStudy (anonymous):

Well I mean the central angle would be from the center, and there are 360 degrees in a circle. If 47% of it is taken up by the section mentioned in the question, we could just do this: .47*360=169.2, and then rounding to the nearest degree would get you 169 degrees. I think it's how to do it. I could be wrong. I wasn't exactly a fan of geometry when I took it.

OpenStudy (taylovesmatthew):

@halorazer well that was easy lol thank you i have never heard any question like that.

OpenStudy (anonymous):

47% means 47% of area. sector area = (central angle/360)*area So (sector area/area)*100 = (central angle/360)*100 47 = (central angle/360)*100 Now solve for central angle.

OpenStudy (anonymous):

It wasn't given area, sir. No measurements were given aside from the 47% part.

OpenStudy (anonymous):

Do tell how we are supposed to find area.

OpenStudy (anonymous):

Plus, it's referring to a graph, not an actual object. How can you measure the area of somethign that isn't even real? O.o

OpenStudy (anonymous):

Yes it's a graph. You can take any character of circle sector which depends on the central angle. For e.g arc length. But usually, it is taken as area.